# St. Aldates, Oxford, OX1 1DW, UK

### v1

@@ Abstract
We find a dual description of the supersymmetric $SL(2,\mathbb{C})/SU(2)$ WZW model as a super-winding condensate CFT, which follows simply from the bosonic AdS3 duality found in ER_EPR. This duality can be seen as a dimensional uplift of the mirror symmetry between the $H_{3}^{+}/U(1)$ supercoset and $\mathcal{N}=2$ super-Liouville theory.
@@ 1 Introduction
In this note, we find the supersymmetric analogue of the bosonic (E)AdS3 duality found in ER_EPR; Halder_2023. On one side of the bosonic AdS3 duality is the $SL(2,\mathbb{C})/SU(2)\simeq H_{3}^{+}$ WZW model, which is the worldsheet theory of a bosonic string in EAdS3. The dual is a winding condensate CFT given by a linear dilaton + $\beta\gamma$ system deformed by $(1,1)$ winding operators
@@
The AdS3 duality is a dimensional uplift of the FZZ duality, which relates the $H_{3}^{+}/U(1)$ coset CFT to sine-Liouville theory Kazakov_2002. The dualities are related as follows: both the $H_{3}^{+}$ WZW model and the winding condensate CFT have a current $J_{3}$ which represents a timelike isometry within spacetime. Gauging this isometry in the WZW model gives the coset CFT; gauging it in the winding condensate gives sine-Liouville theory.
The FZZ duality has an $\mathcal{N}=2$ supersymmetric analogue, which relates the $H_{3}^{+}/U(1)$ Kazama-Suzuki supercoset theory to $\mathcal{N}=2$ super-Liouville theory Giveon_1999; SUSY_FZZ. In this case, the two theories are related by mirror symmetry SUSY_FZZ. It is natural to seek a corresponding supersymmetric analogue of the AdS3 duality, which would relate the supersymmetric $H_{3}^{+}$ WZW model to some super-winding condensate CFT. [^1] Note that the $H_{3}^{+}$ super-WZW model only has $\mathcal{N}=1$ supersymmetry. It is enhanced to $\mathcal{N}=2$ upon gauging $U(1)$ KAZAMA1989232.
We will show that the supersymmetric AdS3 duality follows trivially from the bosonic case due to the following fact: the supersymmetric $H_{3}^{+}$ WZW model at level $k$ factorizes into a bosonic WZW model at level $k_{B}=k+2$ and three free fermions. We can therefore simply replace the bosonic part of the super-WZW model with its dual winding condensate CFT whilst leaving the fermions unchanged. Hence, the dual to the supersymmetric theory is obtained by simply adding three free fermions to the bosonic winding condensate CFT! The explicit map between normalizable vertex operators and the equivalence of their two- and three-point correlators follows immediately from Halder_2023 due to the factorization of the theory. To verify that this is indeed the intended dual to the super-WZW model, in the sense that it is a ‘dimensional uplift’ of the supersymmetric FZZ duality, we show that this super-winding condensate CFT reduces to $\mathcal{N}=2$ Liouville theory upon gauging the timelike isometry $J_{3}$, in analogy with the bosonic case.
@@ 2.1 The supersymmetric EAdS3 WZW Model
One can decouple the fermions within the SUSY WZW action via a field redefinition. This leaves a bosonic WZW model into $H_{3}^{+}$ at level $k_{B}=k+2$ along with three free fermions. The adjustment to the level comes from a chiral anomaly in the path integral measure DIVECCHIA1985701; Schnitzer:1988qj. Pairing two of the fermions as $\psi_{\pm}=\frac{1}{\sqrt{2}}(\psi_{1}\pm i\psi_{2}),\tilde{\psi}_{\pm}=\frac{1}{\sqrt{2}}(\tilde{\psi}_{1}\pm i\tilde{\psi}_{2})$, the action takes the form
@@
We will later be interested in the asymptotic form of this action as $r\to\infty$. The fermionic part is unaffected. In order to avoid the divergence of $\sinh^{2}(r)$ in the bosonic part of the action, we can introduce the first-order variables $W:=\xi+i\theta$ and an auxilliary field $\chi$ ER_EPR:
@@
This change of variables gives rise to a dilaton profile $\Phi(r)\sim-\ln(\sinh(r))$ in the quantum theory. One can reobtain the action (2.1) by integrating out $\chi$. If we define the normalized fields [^4] The differing normalization of $\rho$ comes from quantum corrections to the action; see ER_EPR.
@@ 2.1.1 Gauging $U(1)$
In order to compare with the supersymmetric FZZ duality, we must gauge the $U(1)$ timelike isometry of $H_{3}^{+}$, yielding the Kazama-Suzuki supercoset theory into $H_{3}^{+}/U(1)$. Semiclassically, this coset theory is a sigma model into a 2D cigar. To see this, we write the action as a gauged WZW model Nakatsu:1991pu; Schnitzer:1988qj; Witten:1991mk. One must replace $\partial\xi\to\partial\xi+A$ in the bosonic part, replace $\partial\to\partial+A$ in the fermionic part, and set $\psi_{3}=0$ in the action (2.1). In the first-order formalism, we must also shift $\partial\bar{W}\to\partial\bar{W}+A$:
@@
This geometry represents a 2D Euclidean black hole Witten:1991yr; DIJGRAAF1992269. In the asymptotic region $r\to\infty$, it approaches a cylinder with radius $\sqrt{k}$. The geometry ends at $r=0$, which is the tip of the cigar.
@@ 2.2 $\mathcal{N}=2$ super-Liouville Theory
The mirror dual to the Kazama-Suzuki coset theory above is $\mathcal{N}=2$ super-Liouville theory SUSY_FZZ; NAKAYAMA_2004,
@@
$+2\pi\Big(\mu\lambda_{+}\tilde{\lambda}_{+}e^{-\sqrt{k}\rho}e^{-i\sqrt{k}(\phi_{L}-\phi_{R})}+\bar{\mu}\lambda_{-}\tilde{\lambda}_{-}e^{-\sqrt{k}\rho}e^{i\sqrt{k}(\phi_{L}-\phi_{R}}\Big)$ (14)
@@
where $\phi\sim\phi+2\pi\sqrt{k}$. We also have a linear dilaton $\Phi(\rho)=-\rho/\sqrt{k}$. The last term is a contact term whose role is to regulate certain divergences. Amplitudes can be obtained by dropping this term and analytically continuing in momentum GREEN1988559; Hosomichi_2006.
@@
$V_{L}=2\pi\left(\mu\lambda_{+}\tilde{\lambda}_{+}e^{-\sqrt{k}\rho}e^{-i\sqrt{k}\phi}+\bar{\mu}\lambda_{-}\tilde{\lambda}_{-}e^{-\sqrt{k}\rho}e^{i\sqrt{k}\phi}\right).$ (17)
As we have seen, both the Kazama-Suzuki supercoset and $\mathcal{N}=2$ Liouville theory asymptotically approach free cylindrical theories with a linear dilaton $\Phi\sim-\rho/\sqrt{k}$ which grows as we probe further into the geometry $\rho\to$-$\infty$. In the supercoset, the strong coupling region which would exist at $\rho=$-$\infty$ is cut off due to the geometry, which ends at the cigar tip $\rho=0$. On the other hand, in the Liouville theory, the geometry is still an infinite cylinder $\rho\in(-\infty,\infty)$, but the potential $V_{L}$ exponentially suppresses the propagation of strings into the strong coupling region. The statement of the supersymmetric FZZ duality is that these CFTs are exactly equivalent to each other. It can be proven using mirror symmetry, as was done in SUSY_FZZ.
@@ 3 Uplifting the duality
The bosonic version of this duality has already been found in ER_EPR; Halder_2023. For the level $k_{B}$ bosonic WZW model, the dual potential $W=\mu W_{+}+\bar{\mu}W_{-}$ is given by
@@
Winding operators of the cylindrical $\beta\gamma$ system with periodicity $\gamma\sim\gamma+2\pi iR$ take the form ER_EPR; Halder_2023; Frenkel:2005ku
@@
We will be interested in the operators with $w=\pm 1$. By themselves, these operators are not states of the coset. They have a nonzero OPE with $j_{3}^{B}$, where $j_{3}^{B}$ is the purely bosonic current $j_{3}^{B}=\sqrt{k+2}\left(\beta+\frac{1}{4}\partial\gamma\right)$ ER_EPR:
@@
The OPE can be canceled by appending the operator $e^{\pm\frac{\sqrt{k+2}}{4}(\gamma+\bar{\gamma})}$, leaving a primary with conformal weights $h=\bar{h}=(k+2)/4$; moreover, the winding OPE (23) is unchanged. To obtain a primary with weight $(1,1)$, we add the term $e^{-\sqrt{k}\rho}$, giving us the final operators ER_EPR; Halder_2023
@@
$T_{F}(z)=i\sqrt{2}\Bigg[$ $\psi_{3}\left(-\sqrt{1+\frac{2}{k}}(\beta+\frac{1}{4}\partial\gamma)-i\sqrt{\frac{2}{k}}\mathopen{:}\psi_{+}\psi_{-}\mathclose{:}\right)$
$+\frac{1}{\sqrt{2}}\mathopen{:}\psi_{+}e^{\gamma/\sqrt{k+2}}\mathclose{:}\left(\partial\rho+\sqrt{1+\frac{2}{k}}(-\beta+\frac{1}{4}\partial\gamma)-i\sqrt{\frac{2}{k}}\mathopen{:}\psi_{+}\psi_{-}\mathclose{:}\right)$
$+\frac{1}{\sqrt{2}}\mathopen{:}\psi_{-}e^{-\gamma/\sqrt{k+2}}\mathclose{:}\left(\partial\rho-\sqrt{1+\frac{2}{k}}(-\beta+\frac{1}{4}\partial\gamma)+i\sqrt{\frac{2}{k}}\mathopen{:}\psi_{+}\psi_{-}\mathclose{:}\right)$
@@
To see how $\mathcal{V}$ reduces to the Liouville operator upon gauging, we can follow the strategy of Giveon_2003; Maldacena:2005hi. We introduce a scalar $X_{3}$ to bosonize the current $J_{3}$,
@@
$\mathcal{V}_{\pm}=e^{\pm\int\partial Ydz^{\prime}+\bar{\partial}\tilde{Y}d\bar{z}^{\prime}}=e^{\pm(Y+\tilde{Y})}$. (35)
@@ 4 Conclusion
In this short note, we have shown that the supersymmetric version of the AdS3 duality of ER_EPR; Halder_2023 follows trivially from the bosonic case due to the decoupling of the fermionic sector in the $H_{3}^{+}$ super-WZW model. To verify this, we have shown that the winding potential of the dual CFT reduces to the $\mathcal{N}=2$ Liouville potential upon gauging the timelike isometry $J_{3}$.
Due to the somewhat unusual decoupling of the fermions in the supersymmetric theory, one could potentially evaluate the partition function of the dual on various manifolds by exploiting the localization arguments of murthy2025localizationstringsgroupmanifolds. The partition function of $H_{3}^{+}$ diverges, so one may need to analytically continue or quotient the CFT first.
It would be interesting to find a proof of this duality using a gauged linear sigma model, as was done for the original supersymmetric FZZ duality in SUSY_FZZ. Such a linear sigma model may provide insight into the black hole/string transition in AdS3 chen2021blackholestringtransition; Urbach_2022; urbach2023blackholestringtransitionads3; agia2023ads3stringstarspure.

### v2

@@ Abstract
We find a dual description of the supersymmetric $SL(2,\mathbb{C})/SU(2)$ WZW model as a super-winding condensate CFT, which follows simply from the bosonic AdS3 duality found in [arXiv:2104.07233]. This duality can be seen as a dimensional uplift of the mirror symmetry between the $H_{3}^{+}/U(1)$ supercoset and $\mathcal{N}=2$ super-Liouville theory.
@@ 1 Introduction
In this note, we find the supersymmetric analogue of the bosonic (E)AdS3 duality found in Jafferis and Schneider (2022); Halder et al. (2023). On one side of the bosonic AdS3 duality is the $SL(2,\mathbb{C})/SU(2)\simeq H_{3}^{+}$ WZW model, which is the worldsheet theory of a bosonic string in EAdS3. The dual is a winding condensate CFT given by a linear dilaton + $\beta\gamma$ system deformed by $(1,1)$ winding operators
@@
The AdS3 duality is a dimensional uplift of the FZZ duality, which relates the $H_{3}^{+}/U(1)$ coset CFT to sine-Liouville theory Kazakov et al. (2002). The dualities are related as follows: both the $H_{3}^{+}$ WZW model and the winding condensate CFT have a current $J_{3}$ which represents a timelike isometry within spacetime. Gauging this isometry in the WZW model gives the coset CFT; gauging it in the winding condensate gives sine-Liouville theory.
The FZZ duality has an $\mathcal{N}=2$ supersymmetric analogue, which relates the $H_{3}^{+}/U(1)$ Kazama-Suzuki supercoset theory to $\mathcal{N}=2$ super-Liouville theory Giveon and Kutasov (1999); Hori and Kapustin (2001). In this case, the two theories are related by mirror symmetry Hori and Kapustin (2001). It is natural to seek a corresponding supersymmetric analogue of the AdS3 duality, which would relate the supersymmetric $H_{3}^{+}$ WZW model to some super-winding condensate CFT. [^1] Note that the $H_{3}^{+}$ super-WZW model only has $\mathcal{N}=1$ supersymmetry. It is enhanced to $\mathcal{N}=2$ upon gauging $U(1)$ Kazama and Suzuki (1989).
We will show that the supersymmetric AdS3 duality follows trivially from the bosonic case due to the following fact: the supersymmetric $H_{3}^{+}$ WZW model at level $k$ factorizes into a bosonic WZW model at level $k_{B}=k+2$ and three free fermions. We can therefore simply replace the bosonic part of the super-WZW model with its dual winding condensate CFT whilst leaving the fermions unchanged. Hence, the dual to the supersymmetric theory is obtained by simply adding three free fermions to the bosonic winding condensate CFT! The explicit map between normalizable vertex operators and the equivalence of their two- and three-point correlators follows immediately from Halder et al. (2023) due to the factorization of the theory. To verify that this is indeed the intended dual to the super-WZW model, in the sense that it is a ‘dimensional uplift’ of the supersymmetric FZZ duality, we show that this super-winding condensate CFT reduces to $\mathcal{N}=2$ Liouville theory upon gauging the timelike isometry $J_{3}$, in analogy with the bosonic case.
@@ 2.1 The supersymmetric EAdS3 WZW Model
One can decouple the fermions within the SUSY WZW action via a field redefinition. This leaves a bosonic WZW model into $H_{3}^{+}$ at level $k_{B}=k+2$ along with three free fermions. The adjustment to the level comes from a chiral anomaly in the path integral measure Di Vecchia et al. (1985); Schnitzer (1989). Pairing two of the fermions as $\psi_{\pm}=\frac{1}{\sqrt{2}}(\psi_{1}\pm i\psi_{2}),\tilde{\psi}_{\pm}=\frac{1}{\sqrt{2}}(\tilde{\psi}_{1}\pm i\tilde{\psi}_{2})$, the action takes the form
@@
We will later be interested in the asymptotic form of this action as $r\to\infty$. The fermionic part is unaffected. In order to avoid the divergence of $\sinh^{2}(r)$ in the bosonic part of the action, we can introduce the first-order variables $W:=\xi+i\theta$ and an auxilliary field $\chi$ Jafferis and Schneider (2022):
@@
This change of variables gives rise to a dilaton profile $\Phi(r)\sim-\ln(\sinh(r))$ in the quantum theory. One can reobtain the action (2.1) by integrating out $\chi$. If we define the normalized fields [^4] The differing normalization of $\rho$ comes from quantum corrections to the action; see Jafferis and Schneider (2022).
@@ 2.1.1 Gauging $U(1)$
In order to compare with the supersymmetric FZZ duality, we must gauge the $U(1)$ timelike isometry of $H_{3}^{+}$, yielding the Kazama-Suzuki supercoset theory into $H_{3}^{+}/U(1)$. Semiclassically, this coset theory is a sigma model into a 2D cigar. To see this, we write the action as a gauged WZW model Nakatsu (1992); Schnitzer (1989); Witten (1992). One must replace $\partial\xi\to\partial\xi+A$ in the bosonic part, replace $\partial\to\partial+A$ in the fermionic part, and set $\psi_{3}=0$ in the action (2.1). In the first-order formalism, we must also shift $\partial\bar{W}\to\partial\bar{W}+A$:
@@
This geometry represents a 2D Euclidean black hole Witten (1991); Dijgraaf et al. (1992). In the asymptotic region $r\to\infty$, it approaches a cylinder with radius $\sqrt{k}$. The geometry ends at $r=0$, which is the tip of the cigar.
@@ 2.2 $\mathcal{N}=2$ super-Liouville Theory
The mirror dual to the Kazama-Suzuki coset theory above is $\mathcal{N}=2$ super-Liouville theory Ivanov and Krivonos (1983); Hori and Kapustin (2001); NAKAYAMA (2004),
@@
$+2\pi\Big(\mu\lambda_{+}\tilde{\lambda}_{+}e^{-\sqrt{k}\rho}e^{-i\sqrt{k}(\phi_{L}-\phi_{R})}+\bar{\mu}\lambda_{-}\tilde{\lambda}_{-}e^{-\sqrt{k}\rho}e^{i\sqrt{k}(\phi_{L}-\phi_{R})}\Big)$
@@
where $\phi\sim\phi+2\pi\sqrt{k}$. We also have a linear dilaton $\Phi(\rho)=-\rho/\sqrt{k}$. The last term is a contact term whose role is to regulate certain divergences. Amplitudes can be obtained by dropping this term and analytically continuing in momentum Green and Seiberg (1988); Hosomichi (2006).
@@
$V_{L}=2\pi\left(\mu\lambda_{+}\tilde{\lambda}_{+}e^{-\sqrt{k}\rho}e^{-i\sqrt{k}(\phi_{L}-\phi_{R})}+\bar{\mu}\lambda_{-}\tilde{\lambda}_{-}e^{-\sqrt{k}\rho}e^{i\sqrt{k}(\phi_{L}-\phi_{R})}\right).$ (16)
As we have seen, both the Kazama-Suzuki supercoset and $\mathcal{N}=2$ Liouville theory asymptotically approach free cylindrical theories with a linear dilaton $\Phi\sim-\rho/\sqrt{k}$ which grows as we probe further into the geometry $\rho\to$-$\infty$. In the supercoset, the strong coupling region which would exist at $\rho=$-$\infty$ is cut off due to the geometry, which ends at the cigar tip $\rho=0$. On the other hand, in the Liouville theory, the geometry is still an infinite cylinder $\rho\in(-\infty,\infty)$, but the potential $V_{L}$ exponentially suppresses the propagation of strings into the strong coupling region. The statement of the supersymmetric FZZ duality is that these CFTs are exactly equivalent to each other. It can be proven using mirror symmetry, as was done in Hori and Kapustin (2001).
@@ 3 Uplifting the duality
The bosonic version of this duality has already been found in Jafferis and Schneider (2022); Halder et al. (2023). For the level $k_{B}$ bosonic WZW model, the dual potential $W=\mu W_{+}+\bar{\mu}W_{-}$ is given by
@@
Winding operators of the cylindrical $\beta\gamma$ system with periodicity $\gamma\sim\gamma+2\pi iR$ take the form Jafferis and Schneider (2022); Halder et al. (2023); Frenkel and Losev (2006)
@@
We will be interested in the operators with $w=\pm 1$. By themselves, these operators are not states of the coset. They have a nonzero OPE with $j_{3}^{B}$, where $j_{3}^{B}$ is the purely bosonic current $j_{3}^{B}=\sqrt{k+2}\left(\beta+\frac{1}{4}\partial\gamma\right)$ Jafferis and Schneider (2022):
@@
The OPE can be canceled by appending the operator $e^{\pm\frac{\sqrt{k+2}}{4}(\gamma+\bar{\gamma})}$, leaving a primary with conformal weights $h=\bar{h}=(k+2)/4$; moreover, the winding OPE (22) is unchanged. To obtain a primary with weight $(1,1)$, we add the term $e^{-\sqrt{k}\rho}$, giving us the final operators Jafferis and Schneider (2022); Halder et al. (2023)
@@
$T_{F}(z)=i\sqrt{2}\Bigg[$ $\psi_{3}\left(-\sqrt{1+\frac{2}{k}}(\beta+\frac{1}{4}\partial\gamma)-\sqrt{\frac{1}{k}}\mathopen{:}\psi_{+}\psi_{-}\mathclose{:}\right)$
$+\frac{1}{\sqrt{2}}\mathopen{:}\psi_{+}e^{\gamma/\sqrt{k+2}}\mathclose{:}\left(\partial\rho+\sqrt{1+\frac{2}{k}}(-\beta+\frac{1}{4}\partial\gamma)-\sqrt{\frac{1}{k}}\mathopen{:}\psi_{+}\psi_{-}\mathclose{:}-\frac{\partial\gamma}{\sqrt{k(k+2)}}\right)$
$+\frac{1}{\sqrt{2}}\mathopen{:}\psi_{-}e^{-\gamma/\sqrt{k+2}}\mathclose{:}\left(\partial\rho-\sqrt{1+\frac{2}{k}}(-\beta+\frac{1}{4}\partial\gamma)+\sqrt{\frac{1}{k}}\mathopen{:}\psi_{+}\psi_{-}\mathclose{:}+\frac{\partial\gamma}{\sqrt{k(k+2)}}\right)$
@@
To see how $\mathcal{V}$ reduces to the Liouville operator upon gauging, we can follow the strategy of Giveon et al. (2003); Maldacena (2005). We introduce a scalar $X_{3}$ to bosonize the current $J_{3}$,
@@
$\mathcal{V}_{\pm}=e^{-\sqrt{k}\rho}e^{\pm\int\partial Ydz^{\prime}+\bar{\partial}\tilde{Y}d\bar{z}^{\prime}}=e^{-\sqrt{k}\rho}e^{\pm(Y+\tilde{Y})}$. (34)
@@ 4 Conclusion
In this short note, we have shown that the supersymmetric version of the AdS3 duality of Jafferis and Schneider (2022); Halder et al. (2023) follows trivially from the bosonic case due to the decoupling of the fermionic sector in the $H_{3}^{+}$ super-WZW model. To verify this, we have shown that the winding potential of the dual CFT reduces to the $\mathcal{N}=2$ Liouville potential upon gauging the timelike isometry $J_{3}$.
Due to the somewhat unusual decoupling of the fermions in the supersymmetric theory, one could potentially evaluate the partition function of the dual on various manifolds by exploiting the localization arguments of Murthy and Witten (2025). The partition function of $H_{3}^{+}$ diverges, so one may need to analytically continue or quotient the CFT first.
It would be interesting to find a proof of this duality using a gauged linear sigma model, as was done for the original supersymmetric FZZ duality in Hori and Kapustin (2001). Such a linear sigma model may provide insight into the black hole/string transition in AdS3 Chen et al. (2023); Urbach (2022, 2023); Agia and Jafferis (2023).